# Into the Unknown: Self-Learning Large Language Models

### v2

@@ Abstract
We address the main problem of self-learning LLM: the question of what to learn. We propose a self-learning LLM framework that enables an LLM to independently learn previously unknown knowledge through self-assessment of their own hallucinations. Using the hallucination score, we introduce a new concept of Points in the Unknown (PiUs), along with one extrinsic and three intrinsic methods for automatic PiUs identification. It facilitates the creation of a self-learning loop that focuses exclusively on the knowledge gap in Points in the Unknown, resulting in a reduced hallucination score. We also developed evaluation metrics for gauging an LLM’s self-learning capability. Our experiments revealed that 7B-Mistral models that have been finetuned or aligned and RWKV5-Eagle are capable of self-learning considerably well. Our self-learning concept allows more efficient LLM updates and opens new perspectives for knowledge exchange. It may also increase public trust in AI.
Figure 1: The illustrative space of human knowledge embeddings reduced to two dimensions. It visualizes our four methods for the identification of Points in the Unknown (PiU), later exploited in the self-learning loop. Dashed lines are the borders of the Known regions (darker green) – hallucination score thresholds. Out of them are the Unknown regions (lighter green). White points indicate prompts related to knowledge that the model already knows, while red points indicate PiUs. Different shapes depict different methods: (1) circles represent extrinsic (external) triggers, i.e., either user prompts or trending queries; (2) crosses denote open questions-prompts generated by the model itself within a given topic represented by a dotted line; (3) triangles represent the induced questions generated within a topic using 5W+1H; (4) stars indicate the random sampling by selecting random points in the embedding space.
@@ 1 Introduction
Commonly, Large Language Models (LLMs) are pre-trained on large textual corpora and then fine-tuned using additional data to be better adjusted to a given policy or domain. Simultaneously, other quasi-learning methods have been developed, which are based on additional knowledge provided to the model directly in prompts, especially Retrieval Augmented Generation (RAG). In this paper, we explore a different concept: Self-learning LLMs, i.e., persistent acquisition of new knowledge by the model without data provision. Such a process requires several steps: (1) identification of what knowledge to learn, (2) meaningfulness filter to validate whether identified knowledge is worth and reasonable to learn, (3) searching for and gathering new relevant data, and (4) continual model training. Here, we want to break the assumption that the knowledge for fine-tuning must be inflicted beforehand. Simultaneously, we want to exclude cases in which the model learns the knowledge it already knows. For that purpose, we introduce a novel process component: identification of unknown knowledge based on our new concept, Points in the Unknown (PiUs), and placing this in a self-learning loop.
@@
Our contribution presented in this paper covers: (1) definition of Points in the Unknown (PiUs) using the hallucination score; (2) one extrinsic (External Prompt) and three intrinsic (Open and Induced Generation, Oracle-selected) methods to identify PiUs; (3) introduction of Meaningfulness Filter; (4) metrics to gauge the capability of a model to conduct self-learning: Curiosity Score, Knowledge-Limit Awareness Score, and Self-Learning Capability Score; (5) design of the self-learning LLM using methods for identification of PiUs, Meaningfulness Filter, knowledge searching, and model continual learning; (6) experimental validation; (7) software and data publication for reproducibility.
2 Related Work
@@
Hallucination in the context of LLM is the problem of nonsensical or unfaithful information produced by a generative model. Some works have studied the causes of hallucination Kandpal et al. (2023); Lee et al. (2022); Onoe et al. (2022), as well as the detection methods Lin et al. (2022); Min et al. (2023); Azaria and Mitchell (2023); Manakul et al. (2023); Cao et al. (2023); Yin et al. (2023). Some solutions for overcoming hallucination are proposed in Ji et al. (2023b); Luo et al. (2023); Li et al. (2023); Ji et al. (2023a); Liu and Sajda (2023); Liang et al. (2024); Tian et al. (2024). Meanwhile, in Retrieval Augmented Generation (RAG) Lewis et al. (2020), hallucination is avoided by supplying the prompt with some context retrieved from an outside source, allowing more factual generation without updating the model’s knowledge.
@@
Continual Learning is a training paradigm where the model is subjected to various tasks sequentially; in the context of LLM, the tasks typically comprise domain-specific datasets Jang et al. (2022); Jin et al. (2022); Ke et al. (2023). One challenge is preventing catastrophic forgetting, in which the model loses knowledge from previous tasks McCloskey and Cohen (1989); Ratcliff (1990). Solutions for adding or editing knowledge while avoiding catastrophic forgetting have been proposed Kirkpatrick et al. (2017); Zhu et al. (2020); Sinitsin et al. (2020); Ke et al. (2021); Mitchell et al. (2022).
3 Why Self-Learning is needed
@@
Hallucination is a serious problem that hinders many LLM applications. One of its main reasons is the model’s lack of knowledge on a given topic. This problem is typically overcome by finetuning, i.e., additional learning using new data. Simultaneously, determining what the model already knows and what it does not know yet is very difficult, especially if there is limited information on the model’s past training data. Because of this, finetuning is often a very inefficient process. If finetuning focuses on knowledge already acquired by the model, it does not solve the hallucination problem. Then, we needlessly waste a lot of computing resources by merely repeating known knowledge. Therefore, it is essential to identify the knowledge known and unknown to the model and, in the finetuning process, concentrate only on the unknown.
4 The Concepts of The Known and The Unknown, Point in the Unknown (PiU)
@@
We introduce the concepts of The Known and The Unknown in relation to the knowledge space to define the problem more precisely. Human knowledge is an abstract space formed by vectorial knowledge representations; each point in the space represents an atomic piece of knowledge. The Known refers to an area in the human knowledge space where our LLM does not hallucinate, i.e., it possesses the knowledge related to this region. Then, we can define each point in such an area as Point in the Known (PiK, plural: PiKs). On the other hand, The Unknown refers to an area in the human knowledge space where our LLM hallucinates; each point in such a region is called Point in the Unknown (PiU, plural: PiUs). A PiU represents knowledge that our LLM lacks, which we want the model to identify and acquire.
@@
Alternatively, we could have exploited some other methods for $h(x)$, like SAPLMA Azaria and Mitchell (2023). However, it would require access to the model’s internal features, limiting its feasibility. Therefore, we have remained at the method from Manakul et al. (2023).
5 Methods for Identification of PiUs
@@
Identification of PiUs can be done by evaluating the model’s hallucination scores $h(x)$ on some questions-prompts $x$. These questions, or at least the inspiration for these questions, can come from outside the system, supplied by an external entity. In such cases, the method is considered to have an extrinsic nature. On the other hand, we can create a built-in oracle inside the system that guides automatic question generation in a bicameral-mind manner Jaynes (1990), in which case the method is deemed intrinsic. We propose one extrinsic and three intrinsic methods for PiU identification, Figure 1. All of them differently select or generate candidate points $x$ (question-prompts), which, if tested according to $h(x)\geq LIMIT$, may be identified as PiUs.
5.1 External Prompt (extrinsic)
@@
There are some existing ideas related to collecting prompts that cause hallucination and constructing a dataset based on them for finetuning Zhu et al. (2020); Cao et al. (2023); Tian et al. (2024); Liang et al. (2024). However, they require manual curation of the prompts collected from users or datasets, so the model does not learn fully independently.
@@
In our approach, we utilize an external API to collect trending topics, which are used as inspiration for formulating concrete questions. Every item returned by the API is a list of related phrases; we treat each list as a single topic. Then, the model is asked to generate a specific question $x$ relevant to each topic. Next, the model is asked to answer those questions in order to evaluate its hallucination $h(x)$. Notably, an oracle is still used in the system to control question generation and hallucination scoring, yet the topics that trigger the questions come from outside the system. Unlike the previous ideas, this approach allows the model to continuously learn by itself as long as the external API is available; however, the tested space is limited only to trending topics provided by humans.
5.2 Open Generation (intrinsic)
@@
In this method, the oracle asks the model to propose some topics to learn about. Then, the oracle asks the model to consider those topics and formulate one question ($x$), to which the model believes it does not know the answer. Finally, the oracle asks the model to answer the question $x$ in order to evaluate its hallucination $h(x)$. This method does not require any external entity to work with.
5.3 Induced Generation (intrinsic)
@@
It is based on Five Ws and How, which are widely considered basic questions for information comprehension and data gathering. Here, the oracle also asks the model to propose some topics. Then, the oracle asks the model to formulate a question $x$ using a particular question word; this is repeated six times for what, who, why, where, when, and how, resulting in six different questions. Finally, the oracle also asks the model to answer the questions and evaluate its hallucination $h(x)$. This method also does not require any external inspiration.
5.4 Oracle-Selected (intrinsic)
@@ 6 Self-Learning LLM
Self-Learning is a process where our LLM identifies its own PiUs, searches for the knowledge related to these PiUs, and trains itself on the collected data. It conjures an LLM capable of independent learning to fill the gaps in its own knowledge. Self-learning is made possible by incorporating Self-Questioning – which implements one of the methods for the identification of PiUs – with Knowledge Searching and Model Training in a continuous loop. Self-Learning LLM with an intrinsic method is illustrated in Figure 3, while Self-Learning LLM with an extrinsic method is illustrated in Figure 3.
Self-Learning has some similarities to traditional Continual Learning. The key difference lies in the LLM system, which asks questions by itself and evaluates whether it knows the answer or not. Another difference is the dynamic and integral dataset construction process in Self-Learning; in traditional Continual Learning, the dataset construction process is typically out of the scope. Nevertheless, training techniques and model architectures from traditional continual learning can possibly work in Self-Learning to allow more efficient training and avoid catastrophic forgetting.
6.1 Self-Questioning
Self-Questioning is generally performed through topic generation (or topic collection), question generation, and hallucination scoring. Depending on the selected method, the logical implementation of Self-Questioning may differ. Appendix A provides illustrations of such logical implementations with External Prompt, Open Generation, Induced Generation, and Oracle-Selected.
@@ 6.1 Self-Questioning
Self-Questioning is repeated in a loop for $N$ iterations. The primary output consists of a list of generated questions with hallucination ($Q_{H}$) and a list of generated questions with no hallucination ($Q_{NH}$). In other words, $Q_{H}\subset Unknown$ and $Q_{NH}\subset Known$. We use the NLI-based SelfCheckGPT for hallucination scoring, which takes each sentence in the main passage as a hypothesis and checks each sample as the premise to it, considering only the probabilities of entailment and contradiction to output a normalized score bounded between 0.0 and 1.0 Manakul et al. (2023). A question is categorized into $Q_{H}$ if the average hallucination score of the main passage is greater than 0.5 and into $Q_{NH}$ otherwise.
6.2 Meaningfulness Filter
An optional component before Knowledge Searching is the Meaningfulness Filter. It could be a separately trained model, whose goal is: (1) to remove obviously meaningless PiUs like "How fast cats fly on Mars?"; (2) to implement goals and policy of self-learning, e.g., consider only knowledge (PiUs) relevant to a given domain; (3) to overcome limitations of hallucination score, e.g., when the LLM’s response is "I do not know," and its $h(x)\approx 0$. In our experiments, we did not implement the Meaningfulness Filter because we considered any question in any domain as valid.
6.3 Knowledge Searching
@@
After Self-Questioning and filtering, Knowledge Searching queries an external source to collect knowledge that can answer $Q_{H}$ in order to build the dataset $D_{train}$. Notably, Knowledge Searching may be implemented inside Self-Questioning by immediately searching for the answer to an individual question whenever a hallucination is detected. However, having Knowledge Searching separate is more practical, allowing additional processing without creating a bottleneck for the Self-Questioning process.
6.4 Model Training
@@
The model is trained on $D_{train}$ to absorb the knowledge for answering $Q_{H}$. Once training is done, PiUs should become PiKs, effectively increasing the Known regions and decreasing the Unknown regions of the model. Afterward, the next Self-Learning cycle can start in order to find more PiUs.
7 Metrics for Self-Learning Capability
Self-Learning requires a pretrained model; it can be successfully conducted if the underlying base model already possesses sufficient language understanding and some degree of obedience to instructions. Therefore, we propose some metrics to evaluate the capability of a model to learn independently in a Self-Learning setting.
7.1 Brevity Coefficient
Since the ability to follow instructions is vital for Self-Learning, we use the brevity coefficient to penalize the evaluation when the brevity constraint is violated (e.g., when the model fails to formulate one concrete question without elaboration). It is calculated as follows:
$brev=\begin{cases}1-\frac{\Delta_{len}}{ideal\_len}+\frac{1}{2},&\text{if}\Delta_{len}>50\\ 1,&\text{otherwise}\end{cases}$ (1)
where $ideal\_len$ is the assumed ideal average text length measured in character count, and $\Delta_{len}$ is the average difference between $ideal\_len$ and the texts lengths. More detail is provided in Appendix B.
7.2 Curiosity Score
@@
$Cur=brev*\frac{\#Q_{unique}}{\#Q}$ (2)
where #$\#Q_{unique}$ is the number of unique questions generated, #$\#Q$ is the total number of questions generated, and $brev$ is the brevity coefficient.
#$\#Q_{unique}$ is determined by doing HDBSCAN clustering McInnes et al. (2017) on the question embeddings, counting the number of clusters and outliers. The question embeddings are acquired by using all-MiniLM-L12-v2 from Sentence Transformers Reimers and Gurevych (2019) on the questions generated by our Self-Learning LLM.
7.3 Knowledge-Limit Awareness Score
It indicates how likely a model would come up with a question that it cannot answer without hallucination during Self-Questioning – how likely a model is aware of its own knowledge limitation.

### v3

@@ Abstract
We address the main problem of self-learning LLM: the question of what to learn. We propose a self-learning LLM framework that enables an LLM to independently learn previously unknown knowledge through self-assessment of their own hallucinations. We introduce a concept called Point in the Unknown (PiU) to identify atomic knowledge unknown to a model, along with four methods for automatic PiUs identification, facilitating the creation of a self-learning loop that focuses exclusively on the absorption of currently unknown knowledge into the model. Additionally, we developed evaluation metrics to gauge an LLM’s self-learning capability. Our experiments revealed that LLMs with at least  3B parameters that have undergone some instruction training would be able to perform self-learning well. We further proved the effectiveness of self-learning by comparing the performance of a model that has undergone self-learning to a model that has not. Our self-learning concept allows more efficient LLM updates and opens new perspectives for LLM knowledge exchange.
Index Terms:
self-learning, hallucination, LLM, NLP.
Figure 1: The illustrative space of knowledge embeddings reduced to two dimensions. It visualizes our four methods for the identification of Points in the Unknown (PiUs), later exploited in the self-learning loop. Dashed lines are the borders of the Known regions (darker green) – hallucination score thresholds. Out of them are the Unknown regions (lighter green). White points indicate prompts related to knowledge already known to the model, while red points indicate PiUs. Different shapes depict different methods: (1) circles represent extrinsic (external) triggers, i.e., user queries or trending topics; (2) crosses denote open questions-prompts generated by the model itself within a given topic represented by a dotted line; (3) triangles represent the induced questions generated within a topic using 5W+1H; (4) stars indicate the random sampling by selecting random points in the embedding space.
@@ I Introduction
Commonly, Large Language Models (LLMs) are pre-trained on large textual corpora and then fine-tuned using additional data to be better adjusted to a given policy or domain. Simultaneously, other methods have been developed, which are based on additional knowledge provided to the model directly in the prompt, such as Retrieval Augmented Generation (RAG). In this paper, we explore a different concept: self-learning LLM. It is the persistent acquisition of new knowledge by the model without data provision, taking advantage of three fundamental mechanisms that are integrated in a continuous loop: (1) identification of what knowledge to learn, (2) gathering new relevant data, and (3) continuous model training.
The novelty of this self-learning framework and what sets it apart from traditional continual learning is the ability of the system to determine gaps in its own knowledge and construct the dataset on which it can train itself without repeating knowledge that it already knows. The self-learning LLM utilizes hallucination on simple questions as an indicator of unknown knowledge; while hallucination can be caused by a wide range of factors, one of the main reasons for hallucination on simple questions is due to the model not possessing the factual knowledge that can answer the question [1, 2]. For example, if a model hallucinates when asked, ”Who won the gold medal from men’s singles badminton in the 2024 Summer Olympics?” it is likely that the model’s training did not include the information related to the event.
The self-learning LLM is particularly useful for having an up-to-date model without training a new one from scratch while also minimizing human involvement, greatly reducing development costs. It can be applied in application domains where new facts continuously come up. One example is personalized sentiment prediction and emotion recognition. When new users enter the system, the LLM would be able to perform self-learning only on data related to the new users. Another potential application is the knowledge exchange between LLMs, allowing the knowledge of one LLM to be learned by another LLM automatically.
@@
The contribution presented in this paper covers: (1) the concept of Points in the Unknown (PiUs) to identify knowledge unknown to a model; (2) four different methods to identify PiUs; (3) metrics to gauge the capability of a model to conduct self-learning; (4) design of the self-learning LLM; (5) experimental validation of the methods to identify PiUs; (6) experimental validation of the self-learning LLM; (7) software and data for reproducibility.
II Related Work
@@
Hallucination in the context of LLM is the problem of nonsensical or unfaithful information produced by a generative model. Some works have studied the causes of hallucination [3, 4, 1, 2], as well as the detection methods [5, 6, 7, 8, 9, 10]. Some solutions for overcoming hallucination are proposed in [11, 12, 13, 14, 15, 16, 17, 18, 19]. Meanwhile, in Retrieval Augmented Generation (RAG) [20], hallucination is avoided by supplying the prompt with some context retrieved from an outside source, allowing more factual generation without updating the model.
@@
Continual Learning is a training paradigm where the model is subjected to various tasks sequentially; in the context of LLM, the tasks typically comprise domain-specific datasets [21, 22, 23]. One challenge is preventing catastrophic forgetting, in which the model loses knowledge from previous tasks [24, 25]. Solutions for adding or editing knowledge while avoiding catastrophic forgetting have been proposed [26, 27, 28, 29, 30].
III Why Self-Learning is Needed
@@
Hallucination is a serious problem that hinders many LLM applications. One of its main reasons is the model’s lack of knowledge on a given topic, or the model’s knowledge on the given topic has become obsolete [1, 2]. This problem is typically overcome either by providing the knowledge as additional context [20] without subjecting the model to more learning or by continuous training using new data. With the former, the model would still become outdated after some period of time because most of its knowledge would have become obsolete. With the latter, there is a problem of determining what the model already knows and what it does not know yet, especially if there is limited information on the model’s past training data; if the training focuses on knowledge already acquired by the model, it does not solve the hallucination problem, and we needlessly waste a lot of computing resources by merely repeating known knowledge.
Therefore, it is essential to identify the knowledge known and unknown to the model in order to conduct continuous training as efficiently as possible. The self-learning LLM would be able to distinguish unknown knowledge to create a dataset for its own training automatically. This would reduce the required computing resources as well as human involvement, making continuous training much more efficient. One major concern would be catastrophic forgetting; however, in Section X, we will show that catastrophic forgetting can be avoided by carefully choosing the training technique and architecture.
IV The concept of Point in the Unknown (PiU)
We introduce the concepts of The Known and The Unknown to define the problem more precisely. First of all, if we represent each atomic piece of human knowledge as a vector, it would be possible to form an abstract space that includes all pieces of human knowledge. We call this abstract space as Human Knowledge Space (HKS).
The Known refers to an area in HKS where our LLM does not hallucinate, i.e., it possesses sufficient knowledge related to this region. We call each point in such an area as Point in the Known (PiK, plural: PiKs).
@@ IV The concept of Point in the Unknown (PiU)
The Unknown refers to an area in HKS where our LLM hallucinates; each point in it is called Point in the Unknown (PiU, plural: PiUs). A PiU represents an atomic piece of knowledge that our LLM lacks, which we want the model to identify and acquire.
@@
Alternatively, we could have exploited some other methods for $h(x)$, like SAPLMA [7], DoLa [18], or Lookback Lens [19]. However, SAPLMA and DoLa would require access to the model’s internal features, limiting the feasibility, while Lookback Lens is primarily aimed at RAG applications. Therefore, we decided to stick with SelfCheckGPT, at least for the current work.
SelfCheckGPT has several variants. While the LLM-based variant using GPT-3.5-turbo gave the best results in the original paper, we chose the NLI-based variant. It is recommended if dependency on another LLM is not desired and works faster while still providing decent performance. This variant works by treating each sentence in the main passage as a hypothesis and each sample as a potential premise, using probabilities of entailment and contradiction to output a normalized score bounded between 0.0 and 1.0. In our experiments, we generate 10 samples per main passage, as the original paper’s ablation study showed that performance plateaus at 10 samples, with no significant gain from using more.
V Methods for Identification of PiUs
@@
Identification of PiUs can be done by evaluating the model’s hallucination score $h(x)$ on some question-prompt $x$. The question, or the inspiration for the question, can come from outside the system, supplied by an external entity; such a method would be considered to have an extrinsic nature. On the other hand, we can create a built-in oracle inside the system that guides automatic question generation in a bicameral-mind manner [31], in which case the method is deemed intrinsic. We propose one extrinsic and three intrinsic methods for PiU identification, Figure 1. All of them differently generate $x$, which, if tested according to $h(x)\geq LIMIT$, may be identified as PiU.
V-A External Prompt (extrinsic)
@@
There are some existing ideas related to collecting prompts that cause hallucination and constructing a dataset based on them for finetuning [27, 9, 17, 16]. However, they require manual curation of the prompts collected from users or datasets, so the model does not learn fully independently.
@@
In our approach, we utilize an external API to collect trending topics as inspiration for formulating concrete questions. Every item returned by the API is a list of related phrases; we treat each list as a single topic. Then, the model is asked to generate a specific question $x$ based on each topic. Next, the model is asked to answer the question in order to evaluate its hallucination $h(x)$. This approach allows the model to continuously learn by itself as long as the external API is available; however, the tested space is limited only to trending topics indirectly provided by humans.
V-B Open Generation (intrinsic)
@@
In this method, the oracle asks the model to propose some topics to learn about. Then, the oracle asks the model to consider those topics and formulate one question ($x$), to which the model believes it does not know the answer. Finally, the oracle asks the model to answer the question $x$ in order to evaluate its hallucination $h(x)$.
V-C Induced Generation (intrinsic)
@@
It is based on Five Ws and How, which are widely considered basic questions for information comprehension and data gathering. Here, the oracle also asks the model to propose some topics. Then, the oracle asks the model to formulate a question $x$ using a particular question word; this is repeated six times for what, who, why, where, when, and how, resulting in six different questions. Finally, the oracle also asks the model to answer each question and evaluate its hallucination $h(x)$.
V-D Oracle-Selected (intrinsic)
@@ VI Self-Learning LLM
Self-learning is a process where our LLM identifies its own PiUs, searches for the knowledge related to these PiUs, and trains itself on the collected data. It is made possible by incorporating three fundamental mechanisms in a continuous loop: Self-Questioning, Knowledge Searching, and Model Training. Self-learning LLM with an intrinsic method is illustrated in Figure 3, while self-learning LLM with an extrinsic method is illustrated in Figure 3.
VI-A Self-Questioning
@@
Self-Questioning is generally performed through topic generation (or topic collection), question generation, and hallucination scoring. Depending on the selected method for the identification of PiUs, the logical implementation of Self-Questioning may differ slightly. Self-Questioning is repeated in a loop for $N$ iterations; the primary output consists of a list of generated questions with hallucination ($Q_{H}$) and a list of generated questions with no hallucination ($Q_{NH}$). In other words, $Q_{H}\subset Unknown$ and $Q_{NH}\subset Known$. A question is categorized into $Q_{H}$ if the hallucination score of the main passage is greater than 0.5 or into $Q_{NH}$ otherwise.
VI-B Knowledge Searching
@@
After Self-Questioning and filtering, Knowledge Searching queries an external source to collect knowledge that can answer $Q_{H}$ in order to build the dataset $D_{train}$. Notably, Knowledge Searching may be implemented inside Self-Questioning by immediately searching for the answer to an individual question whenever a hallucination is detected. However, having Knowledge Searching separately is more practical, allowing additional processing without creating a bottleneck for the Self-Questioning process.
VI-C Model Training
@@
The model is trained on $D_{train}$ to absorb the knowledge for answering $Q_{H}$. Once training is done, PiUs should become PiKs, effectively increasing the Known regions and decreasing the Unknown regions of the model. Afterward, the next self-learning cycle can start to find more PiUs.
VII Metrics for choosing a base model for self-learning
Self-learning requires a pretrained model that already possesses a sufficient understanding of instructions. In our initial experiments, we observed that some models are better at asking the ”right” questions (questions on which the model would actually hallucinate) than others. The ability of a model to ask such questions would directly affect the success of self-learning. Therefore, we propose some metrics to evaluate the capability of a model to self-learn.
VII-A Curiosity Score
@@
$Cur=\frac{\#Q_{unique}}{\#Q}$ (1)
where #$\#Q_{unique}$ is the number of unique questions generated, and #$\#Q$ is the total number of questions generated.
#$\#Q_{unique}$ is determined by doing HDBSCAN clustering [32] with the ”leaf” cluster selection method on the question embeddings, counting the number of clusters and outliers. The question embeddings are acquired by using all-MiniLM-L12-v2 from Sentence Transformers [33] on the questions generated by our self-learning LLM.
VII-B Knowledge-Limit Awareness Score
It indicates how likely a model would come up with a question that it cannot answer without hallucination during Self-Questioning – how likely a model is aware of its own knowledge limitation. It is calculated as follows: